how long in minutes does it take to travel 359 feet at 12 mph
Answer by Alan3354(69443)

how long in minutes does it take to travel 359 feet at 12 mph
------------
12 mi/hr = 12 mi/hr * 5280 ft/mi * (1/60) hr/min = 1056 ft/min
----
d = r*t
t = d/r = 359/1056
t =~ 0.33996 minutes